OpenStudy (mollynicole2015):

Vince uses 3 centimeters of tape for every present he wraps. How many presents did Vince wrap if he used 24 centimeters of tape?

OpenStudy (youngstudier):

What do you think you should do first?

OpenStudy (anonymous):

8 presents

OpenStudy (anonymous):

multiply

OpenStudy (mollynicole2015):

Lol I can always count on you. :) 3 * x = 24

OpenStudy (youngstudier):

Exactly, so, we do the reverse on the other side. 3/3 * x = 24 / 3 x = ?

OpenStudy (mollynicole2015):

8

OpenStudy (mollynicole2015):

8 presents

OpenStudy (youngstudier):

Yep. :)

OpenStudy (mollynicole2015):

Thanks Young! ;)

OpenStudy (youngstudier):

You're welcome, Molly :)